3lout:

What is the circumference of the circle? Use 3.14 for π. circle with a segment drawn from the center of the circle to a point on the circle labeled 5 inches 31.40 inches 78.50 inches 15.70 inches 246.49 inches

jhonyy9:

do you know formula ofor circumference of a circle ?

3lout:

yea

3lout:

its (d) C/D = Pi(d)

jhonyy9:



3lout:

idk what that means

jhonyy9:

C circumference C = ? formula

3lout:

C=pi(d)

jhonyy9:

@3lout wrote:
idk what that means
please read the text of your problem ,what is given

jhonyy9:

@3lout wrote:
C=pi(d)
ok but this d what is ?

3lout:

@jhonyy9 wrote:
@3lout wrote:
C=pi(d)
ok but this d what is ?
diameter

jhonyy9:

look please on my posted image

jhonyy9:

@3lout wrote:
@jhonyy9 wrote:
@3lout wrote:
C=pi(d)
ok but this d what is ?
diameter
yes ofc

jhonyy9:

ok but what is this segment with 5 inches ?

jhonyy9:

circle with a segment drawn from the center of the circle to a point on the circle labeled 5 inches what mean this ? do you see this segment on my posted image ?

jhonyy9:

do you know what is a radius ?

jhonyy9:

radius = d/2

3lout:

@jhonyy9 wrote:
circle with a segment drawn from the center of the circle to a point on the circle labeled 5 inches what mean this ? do you see this segment on my posted image ?

1 attachment
3lout:

@jhonyy9 i think that answer is 78.50

KyledaGreat:

The circumference of a circle is calculated by multiplying its diameter by π (pi). To find the diameter, you can use the length of the segment drawn from the center of the circle to a point on the circle labeled 5 inches.

surjithayer:

r=5 d=2r=285=10 \[Circumference=2\pi r=\pi d=3.14*10=31.40~inches\]